OpenStudy (anonymous):

Differential Equations: The definite integral [integral sign from 0 --> infinity] te^[(4-s)t]dt represents the Laplace transform of which function?

OpenStudy (anonymous):

Equation in a better form: \[\int\limits_{0}^{\infty} te ^{(4-s)t}dt\]

OpenStudy (anonymous):

The integrand for the transform is e^(-st)*f(t). So we need to isolate the e^(-st) te^(4-s)t=te^(4t)*e^(-st) by the properties of exponents So we can see f(t) there as te^(4t)

OpenStudy (anonymous):

so that's it? whatever is left over after the e^(-st) was isolated is what was transformed?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

awesome, thanks!